may i know how much of the volume of gas in liter if my balloon radius is 120cm.
Found 2 solutions by rothauserc, Alan3354:
Answer by rothauserc(4718)

Volume(V) of the balloon is (4/3) * pi * (120)^3 = 7238229.4738 cm^3
:
1 cm^3 is 0.001 liters, therefore
:
V in liters = 7238229.4738 * 0.001 = 7238.229 liters of gas
:

Answer by Alan3354(69443)

may i know how much of the volume of gas in liter if my balloon radius is 120cm.
============
If the balloon is a sphere, use the formula for the volume of a sphere.
If you don't know that, learn it and remember it.